Citric acid, H3C6H5O7, is a triprotic acid. Consider a buffer system comprising H2C6H5O7- and HC6H5O72- ions.
What is the net ionic equation for the reaction that occurs when NaOH is added to a buffer containing H2C6H5O7- and HC6H5O72- ?
Answer #1

Net ionic equations are as following: © H3C6H507 + OH = H2C6H507 + H2O @ Hq C H507 + OH = HCcH5 0 + H20 HC H5 0 + OH = Ca H5